Each day the probability of rain on a tropical island is 7/8. Which expression represents the probability that it will rain on the island n days in the next three days?
A) 8C7(3)^n(3)^8-n
B) 3C3(7/8)^3(1/8)^n
C) 3Cn(7/8)n(1/8)^3-n
D) nC3(7/8)3(1/8)^n

Answer by stanbon(75887)

Each day the probability of rain on a tropical island is 7/8. Which expression represents the probability that it will rain on the island n days in the next three days?
---
Binomial Problem with # of trials = 3 ; p = 7/8; q = 1/8; # of successes = n
-----
Ans: 3Cn(7/8)^n(1/8)^(3-n)
==============================
Cheers,
Stan H.
============================
A) 8C7(3)^n(3)^8-n
B) 3C3(7/8)^3(1/8)^n
C) 3Cn(7/8)n(1/8)^3-n
D) nC3(7/8)3(1/8)^n